OpenStudy (anonymous):

I don't know how to do this. Solve the equation. Identify any extraneous solutions. x= square root x +30

OpenStudy (anonymous):

\[x=\sqrt{x+30}\] \[x^2=x+30\]\[x^2-x-30=0\] solve the quadratic

OpenStudy (anonymous):

How do you solve a quadratic?

OpenStudy (anonymous):

this one factors as \[(x-6)(x+5)=0\] so \(x=6\) or \(x=-5\) then check the original equation you will see that \(6\) works and \(-5\) does not

OpenStudy (anonymous):

so what is the extraneous idk what that means

OpenStudy (anonymous):

it means the one that doesn't work

OpenStudy (anonymous):

6 works because \[6=\sqrt{6+30}\] \[6=\sqrt{36}\]\[6=6\] is true

OpenStudy (anonymous):

oh thanks

OpenStudy (anonymous):

-5 does not work because \[-5\neq \sqrt{-5+30}\] \[-5\neq \sqrt{25}\] \[-5\neq 5\]

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

oh see i didn't know that was how it works